a cyclist travels 32 kilometers during the first two hours of riding and 13 kilometers during the next hour. What was his average?
Answer by nerdybill(7384)

a cyclist travels 32 kilometers during the first two hours of riding and 13 kilometers during the next hour. What was his average?
.
Total distance traveled:
2*32 + 1*13
= 64 + 13
= 77 kilometers
.
Total time traveled:
2 + 1 = 3 hours
.
Average:
77/3 = 25.67 kilometers per hour